Currently, sarah is three times her daughters's age. In 13 years, she will be two times her daughters's age. How old is each now?
Answer by ankor@dixie-net.com(22740)

let s = Sarah's age
let d = daughter's age
:
Write an equation for each statement:
:
Currently, sarah is three times her daughter's age.
s = 3d
:
In 13 years, she will be two times her daughter's age.
s + 13 = 2(d+13)
s + 13 = 2d + 26
s = 2d + 26 - 13
s = 2d + 13
replace s with 3d
3d = 2d + 13
3d - 2d = 13
d = 13 is the daughter's age
I'll let you find Sarah's age
:
:
Check your solutions in the 2nd statement